OpenStudy (anonymous):

how do you factor and solve the equation 50x^4-32=0 ??

OpenStudy (campbell_st):

well start by taking out 2 as a common factor \[2(25x^4 - 16) = 0\] this is now an equation that is reducible to a quadratic let \[x^2 = u\] so you have \[2(25u^2 - 16) = 0\] this is the difference of 2 squares... which can be factorised... are you able to do this...?